List the species present in an aqueous solution of each of the following:

a. HF b. KOH c. NH3 d. HCl e. Ba(OH)2

Given the concentrations of the following solutions provide the pH, pOH, H3O, and OH-

a. 0.0025 M HCl b. 0.0025 M NaOH c. 0.035 M Ba(OH)2 d. 0.0015 M H2SO4
Answer #1

a) HF is a molecular compound. It is a weak electrolyte. It is not completely dissociated into water.

HF (aq) + H2O (l)   H3O + (aq) + F - (aq)

Water undergoes auto ionization as 2 H2O (l)   H3O + (aq) + OH - (aq)

Hence, Species present in the aqueous solution of HF are, HF , H3O  +, F - , H2O and OH -

b) KOH is a strong electrolyte. It dissociates completely into water.

KOH (aq) K + (aq) + OH - (aq)

Water undergoes auto ionization as 2 H2O (l)   H3O + (aq) + OH - (aq)

Hence, Species present in the aqueous solution of KOH are, K + , H3O +,OH - and H2O

c) Ammonia ( NH3)

It is a weak electrolyte. It is not completely dissociated into water.

NH3 (aq) + H2O (l)   NH4+ (aq) + OH - (aq)

Water undergoes auto ionization as 2 H2O (l)   H3O + (aq) + OH - (aq)

Hence, Species present in the aqueous solution of NH3 are NH3 , H3O +,NH4+ , H2O and OH -

d) HCl

HCl is a strong electrolyte. It dissociates completely into water.

HCl (aq) + H2O (l)    H3O + (aq) + Cl - (aq)

Water undergoes auto ionization as 2 H2O (l)   H3O + (aq) + OH - (aq)

Hence, Species present in the aqueous solution of HCl are Cl - , H3O +,OH - and H2O

e) Ba(OH) 2

Ba(OH) 2 is a strong electrolyte. It dissociates completely into water

Ba(OH) 2 (aq) Ba 2+(aq) + 2 OH - (aq)

Water undergoes auto ionization as 2 H2O (l)   H3O + (aq) + OH - (aq)

Hence, Species present in the aqueous solution of Ba(OH) 2 are Ba 2+ , H3O +,OH - and H2O

a) 0.0025 M HCl

We have, pH = - log [H3O + ]

Therefore, pH = - log 0.0025 = 2.60

We have, pH + pOH =14

pOH = 14 - pH = 14-2.60 = 11.4

We have, [H3O + ] [OH -]= 10 -14

Therefore, [OH -]= 10 -14 / 0.0025 = 4.0 10 -12 M

ANSWER : pH = 2.60 , pOH = 11.4 , [H3O + ] = 0.0025 M , [OH -] = 4.0 10 -12 M

b) 0.0025 M NaOH

We have, pOH = - log [OH -] = - log 0.0025 = 2.60

We have, pH + pOH =14

pH = 14 - pOH = 14-2.60 = 11.4

We have, [H3O + ] [OH -]= 10 -14

Therefore, [H3O + ] = 10 -14 / 0.0025 = 4.0 10 -12 M

ANSWER : pH = 11.4 , pOH = 2.60 , [H3O + ] =4.0 10 -12 M , [OH -] = 0.0025 M

C) 0.035 M Ba(OH) 2

Consider dissociation of Ba(OH) 2 in water. Ba(OH) 2 (aq) Ba 2+ (aq) + 2 OH - (aq)

From above reaction, [OH -] = 2 [Ba(OH) 2 ] = 2 * 0.035 = 0.070 M

We have, pOH = - log [OH -] = - log 0.070 = 1.15

We have, pH + pOH =14

pH = 14 - pOH = 14-1.15= 12.85

We have, [H3O + ] [OH -]= 10 -14

Therefore, [H3O + ] = 10 -14 / 0.070 = 1.43 10 -13 M

ANSWER : pH = 12.85 , pOH = 1.15 , [H3O + ] =1.43 10 -13 M , [OH -] = 0.070 M

D) 0.0015 M H2SO4  

Consider dissociation of H2SO4 in water. H2SO4 (aq) + H2O (l) 2 H3O + (aq) + SO 42- (aq)

From above reaction,  [H3O + ] = 2 [H2SO4] = 2* 0.0015 = 0.003 M

We have, pH = - log [H3O + ]

Therefore, pH = - log 0.003 = 2.52

We have, pH + pOH =14

pOH = 14 - pH = 14-2.52 = 11.48

We have, [H3O + ] [OH -]= 10 -14

Therefore, [OH -]= 10 -14 / 0.003 = 3.33 10 -12 M

ANSWER : pH = 2.52 , pOH = 11.48 , [H3O + ] =0.003 M , [OH -] = 3.33 10 -12 M